Verify that u=2x+a−−−−−√+2y+b−−−−−√ is a solution of the equation ∂x∂u+∂y∂u=u.
1
Expert's answer
2020-08-04T19:24:16-0400

∂x∂u+∂y∂u=u.

Divide throughout by ∂x∂y, we get


"\frac{\partial u}{\partial \:y}+\frac{\partial u}{\partial \:x}= \frac{\partial{u}}{\partial \:x\partial \:y}"


Given u= "\left(2x+a\right)-\sqrt{2y+b}"

Left side:


"\frac{\partial }{\partial \:x}\left(\left(2x+a\right)-\sqrt{2y+b}\right)+\frac{\partial }{\partial \:y}\left(\left(2x+a\right)-\sqrt{2y+b}\right)"


"\frac{\partial }{\partial \:x}\left(\left(2x+a\right)-\sqrt{2y+b}\right)= 2"


"\frac{\partial }{\partial \:y}\left(\left(2x+a\right)-\sqrt{2y+b}\right)= \frac{1}{-\sqrt{2y+b}}"


Hence left side is "2-\frac{1}{\sqrt{2y+b}}"


Right side:


"\frac{\partial }{\partial \:x\partial \:y}\left(\left(2x+a\right)-\sqrt{2y+b}\right)" = "\frac{\partial }{\partial \:y}(2)" =0


Hence both sides are not equal and hence u is not the solution for given equation.